### v1

# CoVoMix: Advancing Zero-Shot Speech Generation for Human-like Multi-talker Conversations

Abstract
Recent advancements in zero-shot text-to-speech (TTS) modeling have led to significant strides in generating high-fidelity and diverse speech. However, dialogue generation, along with achieving human-like naturalness in speech, continues to be a challenge in the field. In this paper, we introduce CoVoMix: Conversational Voice Mixture Generation, a novel model for zero-shot, human-like, multi-speaker, multi-round dialogue speech generation. CoVoMix is capable of first converting dialogue text into multiple streams of discrete tokens, with each token stream representing semantic information for individual talkers. These token streams are then fed into a flow-matching based acoustic model to generate mixed mel-spectrograms. Finally, the speech waveforms are produced using a HiFi-GAN model. Furthermore, we devise a comprehensive set of metrics for measuring the effectiveness of dialogue modeling and generation. Our experimental results show that CoVoMix can generate dialogues that are not only human-like in their naturalness and coherence but also involve multiple talkers engaging in multiple rounds of conversation. These dialogues, generated within a single channel, are characterized by seamless speech transitions, including overlapping speech, and appropriate paralinguistic behaviors such as laughter. Audio samples are available at https://aka.ms/covomix.
Figure 1: The overview of CoVoMix framework, which consists of a multi-stream text-to-semantic model, a conditional flow-matching based acoustic model for mixed mel-spectrogram generation, and a HiFi-GAN based vocoder for waveform production.
1 Introduction
Zero-shot Text-to-speech (TTS) technology aims to create human-like natural speech with voice characteristics prompted by the context. Recent deep learning advancements have significantly improved synthesized speech quality, especially in formal reading scenarios [1; 2; 3]. However, TTS systems still struggle with rendering spontaneous-style speech and managing seamless transitions during conversations—common occurrences in everyday human communication [4; 5]. On the one hand, spontaneous-style speech encompasses phenomena like filled pauses, interjections, repairs, repetitions, and laughter, which lend human-realistic naturalness to spoken language [6]. On the other hand, in natural conversations, speakers intuitively time their speech, determining when to speak and when to yield, resulting in seamless transitions with appropriate overlaps or moments of silence [7; 8]. Considering these conversational features, we summarize three main challenges in generating spontaneous dialogues.
First, the scarcity of high-quality, spontaneous conversational datasets, along with the difficulty in segmenting paralinguistic behaviors, continues to be a significant obstacle in the field. Spontaneous behavior and non-verbal expressions such as laughter, receive insufficient attention in speech synthesis. Existing high-quality datasets for spontaneous and conversational speech are relatively small and involve a limited number of speakers [9]. Identifying and segmenting these paralinguistic features, as highlighted by studies [10; 11], poses difficulties. Models that require pre-alignment necessitate sophisticated manual annotation. Otherwise, the low-quality data can adversely impact performance, particularly in TTS tasks [12].
Second, research on turn-taking mechanisms in multi-speaker dialogues is less explored. In such dialogues, the forthcoming speaker anticipates the end of the current speaker’s turn by analyzing structural and contextual cues, and then begins their speech seamlessly at the anticipated transition point [13]. Speakers tend to adapt the pause length to match other participants [14]. Overlapping speech occurs when one speaker starts talking before another finishes, which can be a sign of enthusiasm or an attempt to take turns [15].
Third, the consistency in multi-round dialogues is not guaranteed in conventional methods. Simply concatenating each utterance to form a dialogue may result in inconsistent speaker characteristics, particularly when the same speaker engages in multi-round dialogue. In addition, the context of the preceding utterance plays an important role in the control of pauses and prosody, and thus influences the naturalness of generated dialogues [16].
To effectively generate human-like dialogue, we propose CoVoMix, named Conversational Voice Mixture Generation, for multi-talker dialogue generation, shown in Figure 1. The main contributions of the paper can be summarized as follows:
  1. 1.
    To the best of our knowledge, it is the first attempt at zero-shot, human-like multi-speaker multi-round dialogue generation model. We propose 1) a simultaneous multi-stream semantic token prediction from a dialogue text; and 2) a multi-talker flow-matching based acoustic model for generating a mixed mono mel-spectrogram given multiple contexts. It is capable of generating single-channel multi-round dialogue containing multiple speakers concurrently, enabling simultaneous timbre cloning of multiple speakers in zero-shot scenarios.
  2. 2.
    The model, trained on conversational speech, transforms continuous speech features into discrete tokens, capturing paralinguistic behaviors. An autoregressive model is utilized to convert dialogue text, delimited by speaker tags, to these discrete token sequences for each speaker. The overlapping segments, specifically the duration of overlap among speakers, are learned from the training data. Consequently, during the inference stage, these overlaps are automatically generated without the need for any special symbols in the input text.
  3. 3.
    We design a variety of evaluation metrics for dialogue generation, and demonstrate that the CoVoMix model is proficient at generating both human-like dialogues and monologues, exhibiting natural speaker turn-taking, realistic vocal burst-like laughter, consistent speech in terms of speaker similarity throughout multiple rounds of dialogue.
  4. 4.
    We employ the Fisher dataset for this study, which was curated for robust speech recognition. Our approach includes a comprehensive strategy for processing this dataset, including both training and evaluation for monologue and dialogue speech. The data processing script, along with the model training codes and evaluation tools, will be publicly available.
2 Related Work
2.1 Text-to-Speech as Language Modeling
With the advancement in neural codecs for speech, many recent studies have explored token-based language modeling for text-to-speech task. VALL-E [2], a zero-shot text-to-speech model, is a text conditioned language model trained on EnCodec tokens [17]. SPEAR-TTS [18] casts TTS as two tasks: text to high-level semantic tokens and semantic tokens to low-level acoustic tokens. Base-TTS [19] is the largest TTS model to date, which is trained on 100K hours data and utilizes discrete codes extracted from a WavLM Self-Supervised Learning (SSL) speech model [20] as speech representation. HuBERT [21], another model within the SSL framework, serves as the foundational architecture for the Generative Spoken Language Modeling (GSLM) family [22; 8]. The GSLM models leverage HuBERT units specifically for the task of speech continuation, distinguishing themselves by operating directly on raw audio signals without the need for any labels or accompanying text.
2.2 Flow Matching
Continuous Normalizing Flows (CNF) [23] is a deterministic framework that is capable of encompassing the probability paths modeled by diffusion processes [24]. It learns the transformation from normal distribution $p_{0}$ to data distribution $p_{1}$, and it parameterizes a time-dependent vector field $v_{t}$ that pushes samples from prior towards the target distribution [25; 26].
Flow Matching (FM) [25] is a simulation-free approach for training CNFs at scale based on regressing vector fields of fixed conditional probability paths. The relationship between the vector field and the flow $\phi$ is defined via an ordinary differential equation (ODE) $d\phi_{t}(y)=v_{t}(\phi_{t}(y))dt,\phi_{0}(y)=y$ [25]. Benefiting from flow matching models, text-to-speech models, such as VoiceFlow [27], MatchaTTS [28] and Voicebox  [26], can generate high-quality speech efficiently.
2.3 Dialogue Generation
With the application of large-scale data and language modeling, generating continuous conversations becomes possible. dGSLM [8] is built on HuBERT semantic token sequence for speech continuation task, which generates two-channel spoken dialogue auto-regressively. However, the textless design prevents dGSLM from controlling the content of the generated speech, resulting in output that is often less intelligible.
SoundStorm [29] is an iterative generative method that converts semantic tokens into acoustic audio tokens. It can perform monologue and dialogue synthesis. However, it synthesizes dialogue sequentially and in reading style, lacking any spontaneous behaviors or instances of overlapping speech.
CHATS [30], similar to dGSLM [8], is an auto-regressive dialogue generation model, which transforms written dialogues into spoken conversations. It is conditioned on the phonetic transcription and can generate meaningful utterances on the speaker side. However, it is unable to control the generation of spontaneous behaviors for both speakers. Its language model is limited to processing a fixed number of dialogue rounds and can only accommodate a predetermined number of speakers. This restriction narrows its potential use cases.
3 Method
3.1 Problem Formulation
Zero-shot speech generation is a task where a model synthesizes speech in a target voice that was not present in its training data. This task requires only a transcript of what is to be spoken and a speech prompt—a brief sample recording of the target voice. It is generally achieved by in-context learning with a dataset of transcribed speech $\{x,y\}$ where ${y}$ and ${x}$ denote speech utterances and their transcriptions, respectively. Zero-shot multi-talker conversational speech synthesis is designed to generate the voices of multiple speakers simultaneously, based on their transcriptions and prompts. Our approach diverges from the traditional method of synthesizing each voice individually and then concatenating them. The goal is to capture the dynamic nature of real conversations, where participants may speak over each other or respond spontaneously with interjections such as laughter.
Our proposed method, CoVoMix, consists of a multi-stream text-to-semantic model, an acoustic model and a vocoder. The text-to-semantic model first generates multi-stream semantic token sequences for each speaker, given the dialogue transcription. Then the acoustic model transforms these semantic sequences into a mixed mel-spectrogram. A vanilla HiFi-GAN vocoder [31] finally synthesizes mono-channel multi-round dialogue from the mel-spectrogram. We utilize a conversational dataset $D=\{x,y\}$ for training, where $y=[y^{1},y^{2}]$ represents a stereo dialogue featuring two speakers, and $x$ corresponds to the text transcription annotated with speaker tags.
3.2 Multi-stream Text-to-Semantic Model
The multi-stream text-to-semantic model is a sequence-to-sequence model based on encoder-decoder architecture. It takes in a text token sequence generated by a BERT text tokenizer  [32], augmented with special tokens denoting speaker transitions and interjections. The output comprises a multi-stream semantic token sequence. For this study, we focus on a dual-stream setup for a dialogue between two speakers. We employ a pre-trained HuBERT model  [^1] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/dgslm/hubert_fisher [8] as a speech tokenizer to extract the clustered discrete HuBERT hidden units as semantic token sequences and process two channels of waveform, separately. If the dialogues are captured in a single-channel recording, it is necessary to perform speaker separation to produce a dual-channel waveform in our approach. The process of semantic token extraction operates at the frame level, with a time shift of 20 milliseconds, resulting in the presence of duplicated tokens within the sequence. We train this model on a paired speech-text dataset with cross-entropy loss, as
$\mathcal{L}_{t2s}=\sum_{c=1}^{C}\sum_{i}\log p(s_{i}^{(c)}|s_{1:i-1}^{(c)};\theta,x)$ (1)
where $s_{i}$ is the $i$th semantic token and $c$ denotes the $c$th speaker. In order to predict two-stream semantic token sequences, we adopt a strategy wherein we divide the semantic embedding into two distinct segments in the final linear layer of the decoder. Each segment corresponds to a different speaker participating in the conversation. This approach enables the model to capture contextual information not only from each individual speaker but also from their interaction. The dynamic exchange between speakers significantly shapes the semantic content, especially in scenarios involving multi-round conversations.
3.3 Acoustic Model
The acoustic model is a flow-matching based transformer encoder, which generates a mixed mel-spectrogram, given multi-stream semantic token sequences and multi-speaker prompts.
At each timestamp $t\in[0,1]$, a lookup table first embeds the semantic token sequence $s=[s^{1},s^{2}]$ into $s_{emb}=[s_{emb}^{1},s_{emb}^{2}]$ for two speakers. We extract the corresponding mixed mel-spectrogram $m$ and individual mel-spectrogram $[m^{1},m^{2}]$ for each speaker of dialogue $y$. We randomly choose a mask. The masked part $\tilde{m}=m\odot mask$ is to predict, while the seen part $m_{ctx}=[m^{1}\odot(1-mask),m^{2}\odot(1-mask)]$ is considered as prompt.
At each flow step $t$, we sample $w=(1-(1-\sigma_{min})t)\tilde{m_{0}}+tm$)], where $\sigma_{min}$ is a hyper-parameter controlling deviation, $\tilde{m_{0}}$ is sampled from $\mathcal{N}(m|0,\mathbf{I})$. Then, the sample $w$ at flow step $t$, the acoustic prompt $m_{ctx}$, and semantic embedding sequences $s_{emb}$ are concatenated frame-by-frame to obtain an input matrix $W_{input}$. Conditional Flow Matching (CFM) is a per-example training objective, which provides equivalent gradients and does not require explicit knowledge of the intractable target vector field. Therefore, we train the acoustic model to learn the mixed mel-spectrogram with objective as in Eq.2, with $v_{t}(w,m_{ctx},s_{emb};\theta)$ the transformer output with flow $w$ at step $t$.
$\mathcal{L}_{CFM}=\mathbb{E}_{t,q(m,s),p_{0}(m_{0})}\|mask\odot((m-(1-\sigma_{min})\tilde{m_{0}})-v_{t}(w,m_{ctx},s_{emb};\theta))\|^{2}$ (2)
During inference, to sample mixed mel-spectrogram $m$ from learned distribution $p_{1}(m|s,m_{ctx})$, we sample a gaussian noise $m_{0}$ from $p_{0}=\mathcal{N}(m|0,\mathbf{I})$ use an ODE solver to evaluate the flow $\phi_{1}(m_{0})$ given $d\phi_{t}(m_{0})/dt=v_{t}(w,m_{ctx},s_{emb};\theta)$ and $\phi_{0}(m_{0})=m_{0}$.
We also use classifier-free guidance during training for flow-matching model, which is to trade off mode coverage and sample fidelity [26; 33]. During training, the acoustic prompt $m_{ctx}$ and semantic sequences $s_{emb}$ are dropped with $p_{uncond}$. During inference, we use the modified vector field $\tilde{v}_{t}(w,m_{ctx},s_{emb};\theta)$ written in Equation 3 to replace $v_{t}(w,m_{ctx},s_{emb};\theta)$, where $\alpha$ is a hyperparameter controlling the strength of guidance.
$\tilde{v}_{t}(w,m_{ctx},s_{emb};\theta)=(1+\alpha)v_{t}(w,m_{ctx},s_{emb};\theta)-\alpha\tilde{v}_{t}(w;\theta)$ (3)
4 Experimental Setup
4.1 Data Preparation
The dataset used in this work is Fisher Dataset [34], which is a telephone conversation dataset with 2,000h English conversations about various topics. Each dialogue was recorded in two channels with an 8kHz sample rate and an average duration of 10 minutes. We randomly divide the Fisher dataset into train/valid/test sets with 97/1/2/ split. Each set has different speakers.
The data preparation is different for monologue and dialogue. For monologue, following Nemo [35] script [^2] https://gitlab.nrp-nautilus.io/ar-noc/nemo/-/blob/master/scripts/process_fisher_data.py, we slice long dialogues into smaller mono-channel samples, and concatenate them to match minimum duration. We also prepare the corresponding transcripts, extract the corresponding mel spectrogram and semantic token sequence for each sample. For dialogue, we slice long dialogues into shorter, stereo-channel dialogues containing at least two utterances from distinct speakers. We ensure that the first and last sentences of each processed dialogue do not overlap with other dialogues, thus avoiding any extraneous content in the transcriptions. We organize the multi-round dialogue transcript chronologically by the start time of each utterance, delimited by the ‘[spkchange]’ token. The HuBERT speech tokenizer is employed to extract the semantic tokens for each channel. Additionally, we extract the mel-spectrogram from the mixed waveform. The detailed algorithm for dialogue data preparation is described in Appendix C.
4.2 Model Configarations
We develop two text-to-semantic models, named CoSingle and CoMix, and two acoustic models, named VoSingle and VoMix. CoSingle and VoSingle are trained exclusively on monologue data, VoMix is trained on dialogue data, and CoMix is trained on a combination of monologue and dialogue data. In addition, the vanilla HiFi-GAN vocoder is trained on monologue data.
The text-to-semantic model is a transformer-based model with rotary embedding [36]. The encoder has 4 layers and the decoder has 4 layers. We set the dimension of text encoder and CoSingle decoder to 512, and set CoMix decoder to 1024. In order to process multi-stream for multiple talkers, CoMix applies multiple heads for generating semantic token sequences. The acoustic model is based on transformer encoder with rotary embeddings [36] and adaptive RMSNorm [37] for time conditioning, which has 8 layers and hidden dimension of 1024. VoMix and VoSingle have the same architecture except for the first input linear layer. More details of model architecture is demonstrated in Appendix A.
To demonstrate the performance of our methods, the baseline that we compare with is a flow-matching speech synthesis model with phoneme representation, similar to VoiceBox [26] [^3] https://github.com/lucidrains/voicebox-pytorch.git. The baseline contains two models: the acoustic model and the duration model. The acoustic model of baseline is the same as our VoSingle acoustic model, but generates mel-spectrogram from the phoneme sequence. The duration model of baseline is to predict the duration of each phoneme, which is also trained with flow matching objective and has the same architecture with 2 layers and hidden size of 1024.
We train all models from scratch and perform inference on the best performed model on validation set. The text-to-semantic model is trained for 10 epochs, the acoustic model and duration model are trained for 100 epochs. The probability of dropping condition during training is $p_{uncond}=0.3$, and the strength of guidance is $\alpha=0.7$ during inference.
4.3 System Configuration and Evaluation Setting
We built two systems: CoVoSingle and CoVoMix, and evaluate them on both monologue and dialogue testing sets. CoVoSingle contains CoSingle and VoSingle models. CoVoMix system contains CoMix and VoMix. For monologue generation, CoVoSingle and CoVoMix systems directly feed the output of text-to-semantic model into the acoustic model. The acoustic prompt is extracted from another utterance of the target speaker. For dialogue generation, CoVoSingle generate each utterance of the dialogue and concatenate these waveforms according to the order of transcript. CoVoMix receives dialogue transcription as input and synthesizes mono-channel dialogue directly. The acoustic prompts are extracted from another dialogue of target speakers.
4.4 Evaluation Metrics
Objective Metrics: We use cosine speaker similarity (SIM), word error rate (WER), Mel cepstral distortion (MCD)  [^4] https://github.com/chenqi008/pymcd, and NISQA  [^5] https://github.com/gabrielmittag/NISQA to evaluate monologue generation results [38]. SIM measures the cosine similarity between speaker embeddings of generated utterance and the acoustic prompt, extracted from WavLM-TDNN [20]. We use Microsoft Speech Recognition API for WER calculation, which measures the correctness and intelligibility. We use an improved MCD metric that adopts the Dynamic Time Warping (DTW) algorithm to find the minimum MCD between two speeches [39]. NISQA [38] measures the speech quality and naturalness of the synthesized speech.
Subjective Metrics: We perform a human evaluation on the generated monologue and dialogue examples. For monologue, we measure naturalness using comparative mean option score (CMOS). We also use the similarity mean option score (SMOS) between the synthesized and prompt speech to measure the speaker similarity. For dialogue, we measure naturalness using comparative mean option score (CMOS). 15 professional linguistic experts provide judges for all subjective evaluations. The instructions of subjective evaluations are provided in Appendix D.
Dialogue Metrics: We assess the naturalness of the generated dialogue speech by 1) Turn-taking Statistics: Utilizing a pre-trained speaker diarization model  [40; 41] [^6] https://github.com/pyannote/pyannote-audio, we measure the duration of inter- and intra-speaker silences, overlapped speech, and active speech. 2) Para-linguistic behaviors: In this study, we only evaluate languter. Employing a laughter detection tool [42] [^7] https://github.com/jrgillick/laughter-detection, we segment instances of laughter and calculate the total count and average duration of laughter events. and 3) Speech consistency: We generate ten dialogues, each with more than five utterances from the target speaker. We randomly select five three-second segments from the target speaker and then compare the cosine similarity of speaker embeddings among these segments.
5 Result and analysis
5.1 Objective and Subjective metrics
Table 1 shows objective evaluation results for monologue and dialogue generation across various systems.
We observe that the systems leveraging our proposed methods, i.e., CoVoSingle and CoVoMix, achieve higher speaker similarity, lower WER and MCD than baseline on monologue evaluation set. The phoneme-based baseline model requires accurate phoneme-level alignment, however, it is difficult to perform forced alignment using conventional alignment tool [43] [^8] https://github.com/MontrealCorpusTools/Montreal-Forced-Aligner, especially for speech with spontaneous behavior and noisy background. These inaccurate annotations interfere with the model training, and make the model performance severely degraded. By replacing phoneme representation with semantic token sequence, we no longer rely on phoneme-level alignment, thus improving model performance.
Unlike monologue, on the dialogue test set, the ground truth dialogue and CoVoMix have high WER. This is because there is a certain amount of overlapped utterances in the dialogue, but the given transcription is sorted in chronological order of the beginning of each sentence, thus resulting in inconsistencies between transcription and speech in the overlapping parts. Generating each utterance individually and splicing it into a long conversation, such as CoVoSingle, does not have such mismatch, and therefore produces a lower WER. Moverover, constrained by the performance of automatic speaker diarization, we don’t measure speaker similairity for dialogue evaluation set.
In terms of speech quality, we observe that the proposed systems can surpass the ground truth on both monologue and dialogue sets. The flow-matching based acoustic model is able to eliminate background noise, and therefore produces cleaner audio than real data. Since CoVoMix can generate overlapped speech, which may result in a slightly lower NISQA, comparing with CoVoSingle.
Table 1: Objective evaluation results for monologue and dialogue generation across various systems.
Type System SIM $\uparrow$ WER $\downarrow$ MCD $\downarrow$ NISQA $\uparrow$
Monologue GroundTruth 0.59 6.10 / 3.03
Baseline 0.42 15.85 9.45 2.93
CoVoSingle 0.46 9.99 5.03 3.04
CoVoMix 0.46 8.95 5.12 3.01
Dialogue GroundTruth / 14.91 / 2.73
CoVoSingle / 11.77 4.77 2.90
CoVoMix / 19.84 4.25 2.87
The subjective evaluations consistently corroborate the findings of the objective metrics.
5.2 Dialogue Metrics
We define four turn-taking activities in a dialogue: 1) intra speaker pause (silence between active speech of the same speaker), 2) inter speaker silence (silence between active speech of different speaker), 3) overlapped segments, and 4) active speech of each speaker [14; 8]. Figure 2 shows the distribution of each these turn-taking events, and the similarity to the ground truth reflects the model’s ability to model turn-taking in a dialogue. Despite CoVoSingle can synthesize high-quality monologue, they show poor performance on dialogue turn taking events, especially for intra-speaker pause, inter-speaker silence, and overlap control. Simply concatenating monologue in utterance level and in semantic sequence level results in low variance in inter- and intra- speaker silence distribution, which makes the generation sounds like machine conversation [8]. Compared with ground truth, CoVoMix shows high simlarity in these turn-taking events, making the generated dialogue more human-realistic.
Figure 2: Distribution of durations of turn-taking events across models. The blue line and the green line represent the median and mean of each event. The more similar to groundtruth, the better.
We compute the frequency and duration of spontaneous laughter behaviors across the conversation test set and compared these data across models to observe how similar these statistics are to ground truth. As shown in Figure 3, we notice that all these proposed models are able to generate laughter with frequency similar to ground truth, which shows precise control in these human-like behaviors. Furthermore, CoVoMix generates dialogue with more similar average duration of each laughter compared with ground truth, while CoVoSingle tend to synthesize shorter laughter.
Figure 3: Comparison of number and duration of laughter among models
We calculate the speaker similarity between any two pairs of different utterances in a long conversation. Figure 4 shows a heatmap of the cosine similarity between different segments using utterance level concatenation such as CoVoSingle, using semantic sequence level concatenation, and no concatenation such as CoVoMix. A lighter color means that the speakers of the two segments are less similar. As can be seen from the color inconsistencies in the figure, the utterance level concatenation indeed causes dissimilarity in speaker characteristics, especially for non-adjacent utterances. Semantic level concatenation diminishes this speaker drift problem by generating the whole long conversation all at once. Generating the entire semantic sequence and dialogue without concatenation is the best, with the highest speech continuity and consistency.
Figure 4: Speech consistency of different concatenation methods for dialogue generation
6 Conclusion, Limitation and Future Work
We introduced the CoVoMix system for human-like monologue and dialogue generation. The system was composed of an auto-regressive text-to-semantic model and a flow-matching based acoustic model, with semantic token sequence as an intermediate representation. A 2k-hour public phone conversational dataset was leveraged in training the two models of CoVoMix. Through both objective and subjective evaluations, CoVoMix not only achieved high naturalness and zero-shot speaker similarity in monologue and dialogue generations but also demonstrated controllability in dialogue turn-taking and the generation of spontaneous behavior.
We have observed instances of words being omitted or duplicated occasionally in synthesized speech. This was primarily attributed to the text-to-semantic model being an auto-regressive model without forced duration alignment. Additionally, the dataset utilized for this study was sampled at 8 kHz with background noise, factors that contribute to the degradation of speech quality. In future work, we aim to enhance the text-to-semantic model by scaling it up with more data or initializing it with a pre-trained model, and employing super-resolution technology to improve the training data fidelity.
Appendix A Model Architecture
Figure 5 compares the generation pipeline among conventional method and our proposed CoVoSingle and CoVoMix methods. Figure 5(a) shows the conventional monologue generation process with phoneme representation. (b) shows our proposed CoVoSingle approach for monologue generation. (c) demonstrates the concatenation method for conventional and CoVoSingle models to generate dialogue. (d) shows the architecture of our proposed CoVoMix model for monologue and dialogue generation.
Figure 5: Comparison of generation pipeline among conventional method and our proposed CoVoSingle and CoVoMix methods.
Figure 6 shows the architecture of text-to-semantic model. We propose two types of text-to-semantic model: CoSingle and CoMix. CoSingle model has single-stream decoder, while CoMix applies multi-stream decoder to generate multiple semantic token sequences for different speakers, as introduced in Section3.2.
Figure 6: Text-to-semantic model
Figure 7 shows the architecture of acoustic model, a flow-matching based transformer encoder. We propose three types of acoustic model: VoSingle, VoMix and VoMix-stereo. VoSingle is a single stream transformer encoder to generate single talker mel-spectrogram. VoMix and VoMix-stereo have the same architecture except for the last linear layers, which generate mono channel mixed mel-spectrogram and multiple single talker mel-spectrograms respectively.
Figure 7: The acoustic model
Appendix B Additional Experiments
We perform detailed ablation experiments for model combination, training data, and model size. We also evaluate the model performance in monologue and dialogue voice conversion.
B.1 Ablation Study on Model Combination
Table 2 illustrates our proposed systems in monologue evaluation set, which are a combination of text-to-semantic model and acoustic model. We abbreviate text-to-semantic as T2S and stereo as S.
For monologue generation, CoVoSingle and CoVoMix systems directly feed the output of text-to-semantic model into the acoustic model. However, CoVoSinx model set the second semantic sequence as all silence tokens. The acoustic prompt is extracted from another utterance of the target speaker.
We first observe that VoMix mono acoustic model has stronger speaker simulation capability than stereo model, because it is more difficult for the latter to predict two long mel-spectrograms simultaneously with independent speaker characteristics. Moreover, we notice that stereo systems show higher NISQA speech quality, indicating that predicting each channel separately causes less distortion in speech quality. Furthermore, when using the same text-to-semantic model, different acoustic models influence WER, indicating that the pronunciation and rhyme also affect the intelligibility of speech. CoVoSinx achieves better speech intelligibility then CoVoSingle due to the use of VoMix.
Table 2: Objective evaluation on monologue for mono and stereo acoustic model
System T2S Acousic Inference SIM WER NISQA
GroundTruth / / / 0.59 6.10 3.03
CoVoSingle CoSingle VoSingle / 0.46 9.99 3.01
CoVoSinx VoMix Mute 2nd seq 0.46 8.78 3.12
CoVoSinx-S VoMix-S Mute 2nd seq 0.36 10.92 3.26
CoVoMix CoMix VoMix / 0.46 8.95 3.01
CoVoMix-S VoMix-S / 0.36 9.46 3.27
B.2 Ablation Study on Data Augmentation for Text-To-Semantic Model
When the total amount of data is deterministic, the diversity of data, such as data duration and the dialogue content, becomes important. Table 3 shows the performance of text-to-semantic model with different data augmentation method on monologue and dialogue generation in terms of correctness. In order for the text-to-semantic model to correctly generate both short and long dialogues, we need to include both short and long sentences in the training speech. Moreover, we can also simulate dialogue given monologues to increase data diversity. which also increase the prediction accuracy of semantic token sequences. After using monologue data of different lengths and synthetic dialog data, the text-to-semantic model achieves the best performance.
Table 3: Ablation study of data augmentation methods
Training Data WER
Real Dialogue Simu Dialogue Short Monologue Long Monologue Monologue Dialogue
$\times$ $\times$ $\times$ 28.50 22.30
$\times$ $\times$ 11.99 21.82
$\times$ 11.22 20.79
10.54 20.87
B.3 Ablation Study on Discrete Semantic Representation and Model Size
Figure 8compares the speaker similarity of oracle phoneme, predicted phoneme using duration predictor, oracle semantic token, and predicted semantic token using text-to-semantic model under the same architecture of acoustic model.
First, we observe that larger acoustic model improves the speaker similarity of the generated speech as the model layer deepens. Second, the similarity using semantic token sequences is higher than using phoneme and even exceeds oracle phoneme representations. This demonstrates the advantages of using semantic token sequences, which not only avoids forced-alignment and improves the word error rate, but also improves the model’s speaker modeling capabilities. Third, for both phoneme and semantic token, the predicted representations are not as good as oracle representations. The duplicated semantic token sequence is more difficult to predict, leading to a bigger gap between oracle and prediction, indicating further improvement space for the performance and accuracy of text-to-semantic model in the future.
Figure 8: Speaker similarity across acoustic model of different size
B.4 Monologue and Dialogue Voice Conversion
In addition to zero shot speech synthesis, our methods can also achieve voice conversion for single person monologues and multi talker conversations. VoSingle performs voice conversion of dialogue by processing each channel individually and then mix them up, while VoMix model achieves voice conversion simultaneously.
Table 4 demonstrates the objective results in monologue and dialogue scenario. We notice that in addition to achieving high speaker similarity, these systems can also achieve high spectral similarity, indicating the strong zero-shot voice conversion capability of our proposed system. Moreover, VoMix achieves comparable MCD results in monologue set, and performs better than VoSingle in dialogue set, demonstrating the superiority of the multi-talker acoustic model VoMix for dialogue modeling.
Table 4: Objective evaluation on voice conversion for monologue and dialogue generation
Type System SIM MCD
Monologue VoSingle 0.44 5.64
VoMix 0.46 5.68
Dialogue VoSingle / 8.28
VoMix / 7.68
Appendix C Data Preparation
Algorithm 1 illustrates the dialogue data preparation pipeline, introduced in Section4.1. We slice long dialogues into smaller stereo-channel dialogue with at least two utterances of different speakers, and set the first and last sentences of the processed dialogue to not contain overlaps with other dialogues so as to ensure no extraneous content in the transcription.
Algorithm 1 Dialogue Data Preparation
0:  Dialogue recordings $y$, corresponding transcriptions $x$.
1:  Segment dialogues into utterances per speaker: $y^{A}$, $y^{B}$ with transcripts $x^{A}$, $x^{B}$ and identity $z$.
2:  Sort $x^{A},x^{B},y^{A},y^{B}$ by start times into sequences $\mathbf{X},\mathbf{Y},\mathbf{Z}$.
3:  Initialize $cache=[]$, $spkcache=[]$, $OutputDialogue=[]$, $OutputTranscript=[]$.
4:  for each $(x_{new},y_{new},z_{new})$ in zip($\mathbf{X},\mathbf{Y},\mathbf{Z}$do
5:     if $cache$ is empty then
6:        Add $(x_{new},y_{new},z_{new})$ to $cache$; add $z_{new}$ to $spkcache$.
7:     else if StartTime$(y_{new})>$ EndTime$(cache[-1])$ AND $|set(spkcache)|>1$ then
8:        Compile dialogue from $cache$.
9:        Compile transcription from $cache$ by start time with speaker change symbol.
10:        Reset $cache$ and $spkcache$.
11:        Add compiled dialogue to $OutputDialogue$, transcription to $OutputTranscript$.
12:     else if EndTime$(cache[-1])-$ StartTime$(cache[0])>$ maxDuration then
13:        Reset $cache$ and $spkcache$.
14:     else
15:        Continue populating $cache$ and $spkcache$.
16:     end if
17:  end for
18:  return  $OutputDialogue$, $OutputTranscript$.
Then, we extract the corresponding mel-spectrogram of each channel and mixed mel-spectrogram of the waveform after mixing. We prepare the multi-round dialogue transcript, which is ordered by start time of each utterance. Two neighboring same-speaker utterances are concatenated directly, while different speakers’ utterances are separated by [spkchange] token, without explicit processing for overlap labelling. A dialogue transcription preparation example is shown in Figure 9. The HuBERT speech tokenizer is applied to extract semantic token sequence for each channel.
Figure 9: Dialogue transcription preparation
Appendix D Subjective Evaluation Instruction
Figure 10, 11 show the subjective evaluation template for monologue generation. Figure 12 shows the subjective evaluation template for dialogue generation.
Figure 10: CMOS evaluation template for monologue generation
Figure 11: SMOS evaluation template for monologue generation
Figure 12: CMOS evaluation template for dialogue generation